OpenStudy (anonymous):

how do you factor (3x-10)/(x+6)

OpenStudy (pinknabastak):

its already factored...

OpenStudy (mathstudent55):

Correct. There's nothing to factor. Each binomial is already completely factored.

OpenStudy (anonymous):

if f(x)=2 what points of x are on the graph of f(x) when f(x)=(x+2)/(x-6) = 2

OpenStudy (mathstudent55):

Solve for x: \( \dfrac{x + 2}{x - 6} = 2 \)

OpenStudy (anonymous):

I do not understand the steps on how you would solve that.

OpenStudy (mathstudent55):

First, multiply both sides by x - 6, to get rid of the denominator.

OpenStudy (anonymous):

got it x=14

OpenStudy (mathstudent55):

\( (x - 6) \dfrac{x + 2}{x - 6} = 2(x - 6)\) \( \cancel{(x - 6)} \dfrac{x + 2}{\cancel{x - 6}} = 2(x - 6)\) \(x + 2 = 2x - 12 \) \(x = 14\) You need one more step. You have to check the solution in the original problem to make sure the solution will not cause a zero in the denominator.